OpenStudy (anonymous):

find the SLOPE OF THE LINE whose equation is 2x+y=5

OpenStudy (anonymous):

You can write this in slope-intercept form first, which is y = mx + b, m being slope. Thus, 2x+ y = 5 becomes y = -2x + 5. So m = -2, and that’s your slope!